OpenStudy (anonymous):

debbie anderson cuts a pice of wood meassuring 4 5/8 inches into two equal pieces. how long is each piece? the length of each piece is_in.

thomaster (thomaster):

\(\large\left( 4+\dfrac{5}{8}\right)*\dfrac{1}{2}\)

thomaster (thomaster):

\(\large 4*\dfrac{1}{2}=2\) \(\large\dfrac{5}{8}*\dfrac{1}{2}=\dfrac{5*1}{8*2}=\dfrac{5}{16}\) So \(2\dfrac{5}{16}\)

OpenStudy (anonymous):

yes